Three species of the phytoseiid genus Euseius are native to California. E. hibisci (Chant) is generally found only within 80 km of the coast in the warmer, drier southern part of the state. It occurs on Persea americana (avocado) foliage throughout this area and on various Citrus species within a few km of the coast. E. tularensis Congdon is common on Citrus spp. (but very rare on Persea) throughout the inland regions of central and southern California. In regions of sympatry, E. hibisci is usually confined to Persea and E. tularensis to Citrus. E. quetzali McMurtry has the largest geographic range, extending from southern California through the northern coastal regions and to higher elevations in the Sierra Nevada foothills and coastal mountain ranges. It is the only California Euseius species collected from native Quercus (oak) foliage and is seldom observed on cultivated hosts.

Host plant specificity appears to be an important factor in Euseius ecology. The experiments reported here were designed to investigate three mechanisms through which these host plant-specific distribution patterns might be maintained.

1)

Do Euseius species exhibit differences in survival and fecundity when reared on different host substrates?

2)

Do Euseius species detect differences between host substrates and respond to them behaviorally?

3)

Is interspecific competition a factor in limiting Euseius species to particular host plants in areas of sympatry?

Host Plant Substrate Effect on Survival and Fecundity of Euseius species.

Eggs from laboratory cultures of each Euseius species were isolated on leaf arenas made from 2 cm discs of washed Quercus spp. Citrus sinensis (navel orange), and P. americana leaves. Pollen of the ice plant (Malephora crocea), the food used to maintain mite cultures, was provided in excess on each arena. Arenas were checked daily and mortality was recorded through the developmental period. Female and male mites surviving to adulthood were paired, and daily oviposition rate and female longevity were measured.

Immature mortality was relatively low on all host plant substrates for all three Euseius species. It ranged from 0-12% on host plant substrates on which the mites are normally collected in the field, and from 15-28% on leaves of host plants on which they don’t usually occur.

Figure 1 illustrates differences in mean total fecundity as a function of host plant substrate. E. quetzali females laid significantly fewer eggs on Persea leaves than on Quercus or Citrus, and fecundity was significantly lowered in E. tularensis on both Persea and Quercus compared to Citrus. Differences in E. hibisci fecundity were not statistically significant.

Host Substrate Preference

To examine behavioral responses to leaf substrates, “leaf choice” arenas were constructed by joining two washed leaves together along matching out edges. The leaves were held in position with masking tape on the upper surface and placed lower surface up on a watersoaked foam pad. Ten females of either E. tularensis or E. hibisci from culture were placed on each of 10 C. sinensis–Persea, 20 C. limon (lemon)–Persea, and 20 C. limon–C. limon arenas. Pollen was added to each leaf. The number of mites on each half of each arena was recorded after approximately 50 hours.

As illustrated in Figure 2, both E. tularensis and E. hibisci demonstrated an apparent capacity to detect differences in host plant substrate. In the C. sinensis– Persea arenas, E. tularensis showed a preference for C. sinensis and E. hibisci for Persea. E. hibisci exhibited a similar preference for Persea in the Persea– C. limon arenas. Surprisingly, E. tularensis showed no significant preference in this test. Neither species exhibited directional preference or a tendency to aggregate in the C. limon–C. limon arenas.

Interspecific Competition

A replacement series test for interspecific mutual interference was modelled after similar studies of interactions between plant species. The following combinations of inseminated, 7-day-old females were placed on each of four leaf arenas: 10 E. hibisci (H), 10 E. tularensis (T), 9H: 1T, 7H: 3T, 5H: 5T, 3H: 7T, and 1H: 9T. After 5 weeks (about 3 generations) of normal maintenance, all adult mites were collected and identified.

If no competition (mutual interference) occurred, the final number of mites expected would be similar for the two-species and the single-species cultures. The initial and final ratios of the two species would also be similar in the absence of competition.

Table 1 shows that in every two-species combination the total number of females observed at the conclusion of the test was lower than expected in the absence of competition. Figure 3 illustrates the mutual reduction in population growth experienced by both species in the two- as compared to one- species cultures. In all but the 9H: 1T combinations, the total number of adults was only about one-half the number in the single species cultures. These results strongly suggest that at least these two Euseius species may compete in the field when they occur sympatrically.

Summary and Conclusions

All three of the interactions examined in this study may be significant in determining the observed patterns of host plant specificity in Euseius species. Mite reproductive capacity is influenced by the leaf substrate. This may be related to leaf sap feeding, which has been demonstrated in E. hibisci and E. tularensis. These species also exhibit a tendency to migrate away from unsuitable and remain on suitable substrates. Finally, in regions where several species occur, host plant range may be further limited by interspecific competitive interactions.

Several studies have shown that these pollenophagous facultative predators have a significant impact on populations of pest arthropods. Studies of the interactions among Euseius species and their host plants reveal the complex nature of Euseius ecology, and should be taken into consideration in biological control efforts.